# A Large Language Model Evolutionary Algorithm for Automatically Generating Metaheuristics

### v2

@@ III-D Mutation, Selection and Feedback
The mutation and selection step in the LLaMEA framework consists mostly of the construction of the feedback prompt to the LLM in order to generate a new solution. Depending on the selection strategy, either the current-best algorithm $a_{b}$ is given back (plus strategy) to the LLM including the score it had on the BBOB suite, or the last generated algorithm $a_{t}$ (comma strategy) is provided to the LLM in the feedback prompt.
@@
  Feedback prompt template $F$     <role and task prompt>  <List of previously generated algorithm names with mean AOCC score>  <selected algorithm to refine (full code) and mean and std AOCC scores>  Either refine or redesign to improve the algorithm.
@@
We ask the LLM to either make a (small) refinement of an algorithm or redesign it completely, where the latter is analogue to a restart or very large mutation rate in an ES optimization run, increasing its exploration behaviour. In the proposed framework we leave this choice to the LLM itself as our task is simply ”Either refine or redesign to improve the algorithm”. In Section V-B, we analyze how the LLM makes this decision over time, in terms of the generated names as well as the code similarity .
@@
Instead of providing an overall score and standard deviation, we can provide the LLM with more performance details of the generated algorithms. In our case, we can provide additional metrics per BBOB function group, potentially giving the LLM information on what kind of functions the solution works well and on which ones it does not work well. This results in mean AOCC values and standard deviations for each of the five function groups, i.e., $(y_{t,1},\ldots,y_{t,5}),(\sigma_{t,1},\ldots,\sigma_{t,5})$. In our experiments we cover the inclusion of such details versus leaving this information out. When we provided more detailed feedback, we provide the mean AOCC and the standard deviation of the AOCC scores for the separable functions, functions with low or moderate conditioning, unimodal functions with high conditioning, multi-modal functions with adequate global structure and functions with weak global structure (the five BBOB function groups). In practise, it would even be possible to provide the full convergence plots, either as csv or image to the multi-modal LLMs; we leave this option for future research.
@@
In the proposed framework we do not keep the entire run history in every iteration (including all codes $(a_{0},\ldots,a_{t})$), as this becomes more and more expensive as the list grows.
However, the inclusion of a larger set of previous attempts by providing the previous solutions in code with their associated scores could in theory be beneficial. The LLM would be able to act as a kind of surrogate model in predicting the next solution. This would translate to evolutionary computation terms, as to keeping an archive of best solutions or the use of machine learning as surrogate models in surrogate assisted optimization [32, 33].
Instead of providing all codes, we keep a condensed list of algorithm names (generated by the LLM) and their respective scores, $((\text{name}(a_{0}),y_{0}),\ldots,(\text{name}(a_{t}),y_{t}))$, to facilitate in-context learning of the LLM [25, 26]. The purpose of keeping this list and providing it to the LLM is two-fold, namely 1) the LLM could learn what kind of algorithms work well and which work less good (by analysing the algorithm names only), and 2) it makes it less likely that the LLM is generating the same algorithm twice.
@@ IV Experimental Setup
To validate the proposed evolution framework for generating and optimizing metaheuristics, we designed a set of experiments to compare different LLMs, different ES strategies and different levels of detail in the provided evaluation feedback. In addition, the best generated algorithms are analysed and compared to several state-of-the-art baselines on the BBOB suite using additional instances and additional dimension settings.
@@ IV-A Large Language Models
In our experiments we have limited the number of LLMs to the ChatGPT family of models, including GPT-3.5-turbo [34], GPT-4-turbo [35] and the recently released GPT-4o [36]. The LLaMEA framework leverages the OpenAI chat completion API call for querying the LLM. For abbreviating LLM-names, we drop the extension ”turbo” in the following.
@@ IV-C Performance Metrics
$\textit{AOCC}(\mathbf{y})=\frac{1}{B}\sum_{i=1}^{B}\left(1-\frac{\min(\max((y_{i}),\textit{lb}),\textit{ub})-\textit{lb}}{\textit{ub}-\textit{lb}}\right)$ (1)
@@
We aggregate the AOCC scores of all benchmark functions by taking the mean. The final mean AOCC is given as feedback to the LLM in the next step, and is used as best-so-far solution in case an improvement was found.
In addition to the mean AOCC score, any runtime or compile errors occurred during validation are also used to give feedback to the LLM. In case of fatal errors (no execution took place), the mean AOCC is set to the lowest possible score, zero.
@@ V Results and Discussion
In Figure 2, the mean best-so-far algorithm evaluation score (AOCC) per configuration (LLM and feedback strategy) is shown. The shaded region denotes the standard deviation over the $5$ independent runs. Different strategies seem to have different effects depending on the model used. For example, the $(1+1)$-selection strategy seems to be beneficial for GPT-4, but is having a deteriorating effect when using GPT-4o. Overall GPT-4 seems to be better suited for the task overall, and GPT-3.5 is clearly a less favorable option.
Figure 2: Mean convergence curves (best-so-far algorithm scores) over the $5$ different runs for each LLM and strategy. Shaded areas denote the standard deviation of the best-so-far. w/Details denotes that we use a feedback mechanism that provides not just the plain average AOCC (and its standard deviation) but also the average AOCC and standard deviation per BBOB function group as feedback to the LLM.
@@ V-A Novelty and diversity
In Figure 3, a word-cloud is shown with all the sub-strings of generated algorithm names and their occurrence frequency visualized. The generated algorithm names were in Python Camel-case style, so it was easy to split the algorithm names in individual parts. In some cases, the algorithm name was just an abbreviation, these abbreviations are kept as words in this analysis. The most used parts in algorithm names are rather generic, such as “evolution”, “adaptive” and “dynamic”, some of the parts directly refer to existing algorithms such as “harmony” and “firework”, and other strings refer to existing strategies such as “gradient”, “local”, “elite” etc. In general when observing the different generated solutions, we see interesting and novel combinations of existing techniques, such as ”surrogate assisted differential evolution combined with covariance matrix adaptation evolution strategies” and ”dynamic firework optimizer with enhanced local search”. A full list of generated algorithms and their names is provided in our Zenodo repository [31].
@@ V-B Mutation Rates
There are a few interesting observations we can make: The difference between parent and offspring for GPT-3.5 is on average much smaller than for other models, the $(1+1)$-GPT3.5-LLaMEA shows much higher differences in the beginning of the run than the other strategies with the same model. GPT-4-LLaMEA in general shows the largest differences (exploration) over the runs. It is very interesting to observe that the ratio of code differences in most cases seems to converge, indicating more exploration in the beginning of the search and more exploitation during the final parts of the search. This is interesting as the LLM has no information on the search budget $T$ of Algorithm 1, and can therefore also not base its decision to make large or small refinements on the stage of the optimization run. We can furthermore see in the generated names of the algorithms that the LLM has either tried to refine the algorithm (adding “Improved” or “Refined” or “Enhanced” to the name, or names such as <algorithm>V1, etc.) or redesign the algorithm using a different strategy (based on different existing algorithm names such as Differential Evolution, Particle Swarm Optimization etc.). We can therefore also look at the similarity between parent and offspring algorithm names to visualize the refinement process of the optimization runs.
@@
$\text{Jaro}(s,t)=\left\{\begin{array}[]{ll}0&\text{if }m=0,\\ \frac{1}{3}\left(\frac{m}{|s|}+\frac{m}{|t|}+\frac{m-t}{m}\right)&\text{otherwise},\end{array}\right.$
@@
Figure 5 illustrates the Jaro similarity mean (five runs) of the subsequently generated algorithm names (i.e., $\text{Jaro}(\text{name}(a^{\prime}),\text{name}(a_{t+1}))$ for $t=0,\ldots,B-1$, where $a^{\prime}=a_{b}$ for $(1+1)$-selection and $a^{\prime}=a_{t}$ for $(1,1)$-selection) over iterations $t$ for each of the nine LLM-based optimization run configurations. In Figure 6, we show the Jaro similarity for just a single run selected for three of the nine configurations.
@@ VI Analysis of Best Algorithms
The experiment above resulted in 3 65736573\,6573 657 algorithms (out of a maximum2225555 runs with $T=100$ iterations, each, for $9$ different models. of 4 50045004\,5004 500) that were at least able to get an AOCC score larger than zero on the BBOB suite of optimization problems in $5$ dimensions. In the next step to validate our proposed framework, we evaluate how much the best of these algorithms can generalize beyond the evaluation performed during the optimization loop, and subsequently we analyze in-depth the code and behaviour of the best-performing algorithm that beats the state-the-art baselines. The evaluation is done in $5$, $10$ and $20$ dimensions, using $5$ different instances and $5$ independent runs per instance (so $25$ runs per BBOB function to optimize). For a fair comparison, we decided to re-run the generated algorithms for $d=5$ in this setting (with $25$ runs), too. The state-of-the-art baselines are the covariance matrix adaptation evolution strategies (CMA-ES) [41], Differential Evolution (DE) [42] and an extensively optimised version of Modular CMA-ES [37] denoted as CMA-best. The optimised CMA-ES algorithm is the best-configured algorithm (in terms of AOCC) on BBOB in $5$ dimensions over more than 52 1285212852\,12852 128 CMA-ES and DE algorithm variants, note that beating this optimized baseline is incredibly hard as it was the result of a very large modular optimization experiment. We use the CMA-best baseline to visualise a reasonable upper-bound of what is possible when optimizing an algorithm configuration to a specific benchmark set and dimensionality.
@@
Figure 7: The empirical attainment functions (EAF) estimate the percentage of runs that attain an arbitrary target value not later than a given runtime. EAF for the best algorithms per LLM configuration, averaged over all 24 BBOB functions in $5d$
In Figure 7 the empirical attainment functions (EAF) [38] are shown for the best algorithm generated per LLM and strategy, so $9$ algorithms in total plus the $3$ baselines. The empirical attainment function estimates the percentage of runs that attain an arbitrary target value not later than a given runtime. Taking the partial integral of the EAF results in a more accurate version of the Empirical Cumulative Distribution Function, since it does not rely on discritisation of the targets. The area under this curve is a similar measure as the AOCC used during the LLaMEA optimization runs, and denotes how well an optimization algorithm performed over the complete run (on average over all instances, independent runs and objective functions). From the EAF curves and the area under these curves (see Table II in the Appendix), we can observe that LLaMEA has found one algorithms (ERADS_QuantumFluxUltraRefined, for short ERADS) with better performance for $d=5$ than the state-of-the-art CMA-ES in terms of AOCC, two algorithms (ERADS and AdaptiveDifferentialEvolutionHarmonySearch) that on average perform better than the optimized CMA-best after the total budget in $5d$ (the EAF curve reaches higher) and six algorithms that perform better than the DE baseline but less than the CMA baseline Detailed convergence curves per BBOB function (for $d=5$) are shown in Figure 8. From this Figure we can observe that especially for BBOB $f_{17}$ and $f_{18}$, the ERADS algorithm shows very promising search behaviour.
Figure 8: Median best-so-far function value ($y$-axis) over the 10 0001000010\,00010 000 function evaluations ($x$-axis) per BBOB function in $5d$ for the best algorithm per LLM configuration (9 algorithms) and the baselines; CMA-ES, DE and CMA-best.
@@
38:return $\mathbf{x}_{opt},f_{opt}$triangleright Return best solution and fitness
The main differences with DE are the use of a memory factor to guide the mutation in certain directions and an adaptive $F$ and $CR$ strategy. It is unclear, why the LLM generated the ”Quantum Flux” component in the algorithm’s name, but it can be assumed that LLM hallucinations [43] will also occur when applied to code generation tasks.
@@
The key differences primarily revolve around the use of a memory vector and a memory factor in combination with parameter adaptation. JADE uses an optional archive with the best solutions which primarily serves to maintain diversity. Furthermore, JADE also features adaptive control parameters (F and CR) that adjust dynamically based on the success rates of previous generations, aiming to optimise these parameters in real time to improve performance. It is interesting to observe that most of the best algorithms proposed by the LLM are similar to Differential Evolution (or Hybrids of CMA and DE), this could indicate a bias towards this kind of algorithms in the LLM, or it could indicate that overall these kind of algorithms work well for the specific benchmark suite.

### v3

@@ III-D Mutation, Selection and Feedback
The mutation and selection step in the LLaMEA framework consists mostly of the construction of the feedback prompt to the LLM in order to generate a new solution. Depending on the selection strategy, either the current-best algorithm $a_{b}$ is given back ($(1+1)$-strategy) to the LLM including the score it had on the BBOB suite, or the last generated algorithm $a_{t}$ ($(1,1)$-strategy) is provided to the LLM in the feedback prompt $F$.
@@
  Feedback prompt template $F$     <Task prompt S>  <List of previously generated algorithm names with mean AOCC score>  <selected algorithm to refine (full code) and mean and std AOCC scores>  Either refine or redesign to improve the algorithm.
@@
We ask the LLM to either make a (small) refinement of an algorithm or redesign it completely, where the latter is analogue to a restart or very large mutation rate in an Evolutionary Algorithm optimization run, increasing its exploration behaviour. In the proposed framework we leave this choice to the LLM itself as our task is simply ”Either refine or redesign to improve the algorithm”. In Section V-B, we analyze how the LLM makes this decision over time, in terms of the generated algorithm names as well as code similarity .
@@
Instead of providing an overall score and standard deviation, we can provide the LLM with more performance details of the generated algorithms. In our case, we can provide additional metrics per BBOB function group, potentially giving the LLM information on what kind of functions the solution works well and on which ones it does not work well. This results in mean AOCC values and standard deviations for each of the five function groups, i.e., $(y_{t,1},\ldots,y_{t,5}),(\sigma_{t,1},\ldots,\sigma_{t,5})$. In our experiments we cover the inclusion of such details versus leaving them out.
@@
In the proposed framework we do not keep the entire run history in every iteration (including all codes $(a_{0},\ldots,a_{t})$), as this becomes more and more expensive as the list grows. However, the inclusion of a larger set of previous attempts by providing the previous solutions in code with their associated scores could in theory be beneficial. The LLM would be able to act as a kind of surrogate model in predicting the next solution. This would translate to evolutionary computation terms, as to keeping an archive of best solutions or the use of machine learning as surrogate models in surrogate assisted optimization [40, 41].
Instead of providing all codes, we keep a condensed list of algorithm names (generated by the LLM) and their respective scores, $((\text{name}(a_{0}),y_{0}),\ldots,(\text{name}(a_{t}),y_{t}))$, to facilitate in-context learning of the LLM [31, 32]. The purpose of keeping this list and providing it to the LLM is two-fold, namely (i) the LLM could learn what kind of algorithms work well and which work less good (by analyzing the algorithm names only), and (ii) it makes it less likely that the LLM is generating the same algorithm twice.
@@ IV Experimental Setup
To validate the proposed evolution framework for generating and optimizing metaheuristics, we designed a set of experiments to compare different LLMs, different ES strategies and different levels of detail in the provided evaluation feedback. In addition, the best generated algorithms are analyzed and compared to several state-of-the-art baselines on the BBOB suite using additional instances and additional dimension settings.
@@ IV-A Large Language Models
In our experiments we have limited the number of LLMs to the ChatGPT family of models, including gpt-3.5-turbo-0125 [42], gpt-4-turbo-2024-04-09 [43] and the recently released gpt-4o-2024-05-13 [44]. The LLaMEA framework leverages the OpenAI chat completion API call for querying the LLM. Each model was run with default parameters (top_p equal to 1) and a temperature of $0.8$. For abbreviating LLM-names, we drop the extension ”turbo” in the following.
@@ IV-C Performance Metrics
$\textit{AOCC}(\mathbf{y}_{a,f})=\frac{1}{B}\sum_{i=1}^{B}\left(1-\frac{\min(\max((y_{i}),\textit{lb}),\textit{ub})-\textit{lb}}{\textit{ub}-\textit{lb}}\right)$ (2)
@@
We aggregate the AOCC scores of all 24 BBOB benchmark functions $f_{1},\ldots,f_{24}$ by taking the mean over functions and their instances, i.e., for an algorithm $a$,
$\textit{AOCC}(a)=\frac{1}{3\cdot 24}\sum_{i=1}^{24}\sum_{j=1}^{3}\textit{AOCC}(\mathbf{y}_{a,f_{ij}})\;.$ (3)
The final mean AOCC over $k=5$ independent runs of algorithm $a$ over all BBOB functions is given as feedback to the LLM in the next step, and is used as best-so-far solution in case an improvement was found. In other words, the quality measure $f(a)$ used in Algorithm 1 is defined as
[…]
In addition to this mean AOCC score, any runtime or compile errors that occurred during validation are also used to give feedback to the LLM. In case of fatal errors (no execution took place), the mean AOCC is set to the lowest possible score, zero.
@@ V Results and Discussion
In Figure 2, the median best-so-far algorithm evaluation score (AOCC) per configuration (LLM and feedback strategy) is shown.
Figure 2: Mean convergence curves (best-so-far algorithm scores) over the $5$ different runs for each LLM and strategy, including the state-of-the-art baseline EoH algorithm. Shaded areas denote the variance of the best-so-far. w/Details denotes that we use a feedback mechanism that provides not just the plain average AOCC (and its standard deviation) but also the average AOCC and standard deviation per BBOB function group as feedback to the LLM.
[…]
From Figure 2 we conclude that different strategies yield a range of different results, depending on the model used. For example, the $(1+1)$-selection strategy seems to be beneficial for GPT-4, but is having a deteriorating effect when using GPT-4o. Overall GPT-4 seems to be better suited for the task overall, and GPT-3.5 is clearly a less favorable option. We also compare to the EoH algorithm with GPT-4 as our baseline. To use EoH, we extended it to work in tandem with IOHexperimenter in the same way as LLaMEA. The $(1+1)$-GPT-4-ES shows better convergence (area under the AOCC curve) than the EoH approach and also results in an overall better algorithm (AOCC) in the end. The EoH approach follows as the second best approach, EoH is run $5$ times with a population size of $5$ and default hyperparameters. EoH performs a bit more stable (in terms of variation over different runs), due to the larger population size. However, it suffers from tries where the generated code is not executable since it has no self-debugging capabilities (unlike LLaMEA). Another limitation of EoH is that it can only deal with single functions, while in our proposed approach complete Python classes are generated, allowing for more complex interactions.
@@ V-A Novelty and diversity
In Figure 3, a word-cloud is shown with all the sub-strings of generated algorithm names and their occurrence frequency visualized. Note that the word-cloud serves purely as an intuitive visual representation to quickly get an overview of the different words the algorithm names contain. All algorithm names and codes are available in our open-source repository [39]. The generated algorithm names are in Python Camel-case style, such that it is easy to split the algorithm names into individual parts. In some cases, the algorithm name is just an abbreviation, and these abbreviations are kept as words in this analysis. The most used parts in algorithm names are rather generic, such as “evolution”, “adaptive” and “dynamic”. Some of the parts directly refer to existing algorithms, such as “harmony” and “firework”, and other strings refer to existing strategies, such as “gradient”, “local”, “elite” etc. In general when observing the different generated solutions, we see interesting and novel combinations of existing techniques, such as ”surrogate assisted differential evolution combined with covariance matrix adaptation evolution strategies” and ”dynamic firework optimizer with enhanced local search”. A full list of generated algorithms and their names is provided in our Zenodo repository [39]. The algorithm names are generated automatically by the LLM that is used by LLaMEA. We have performed manual checks and can confirm that the algorithm names are almost always in line with the actual code. This means that the LLM indeed creates a descriptive name for the generated algorithm, which is based on the key algorithmic components used in the algorithm.
@@ V-B Mutation Rates
There are a few interesting observations we can make: The difference between parent and offspring for GPT-3.5 is on average much smaller than for other models, and the $(1+1)$-GPT3.5-LLaMEA shows much higher differences in the beginning of the run than the other strategies with the same model. GPT-4-LLaMEA in general shows the largest differences (exploration) over the runs. It is very interesting to observe that the ratio of code differences in most cases seems to converge, indicating more exploration in the beginning of the search and more exploitation during the final parts of the search. This is interesting as the LLM has no information on the search budget $T$ of Algorithm 1, and can therefore also not base its decision to make large or small refinements on the stage of the optimization run. When we look at specific code-diffs between parents and offspring generated by the $(1+1)$-GPT-4-LLaMEA, we observe that the LLM mutates both hyperparameter values and higher level logic, like introducing a new crossover or mutation operator. In addition it mutates the comments in the code to reflect and argue about the changes. See the supplemental material for some specific examples.
We can furthermore see in the generated names of the algorithms that the LLM has either tried to refine the algorithm (adding “Improved” or “Refined” or “Enhanced” to the name, or generating names such as <algorithm>V1, etc.) or redesign the algorithm using a different strategy (based on different existing algorithm names such as Differential Evolution, Particle Swarm Optimization etc.). We can therefore also look at the similarity between parent and offspring algorithm names to visualize the refinement process of the optimization runs.
@@
$\text{Jaro}(s,t)=\left\{\begin{array}[]{ll}0&\text{if }m=0\\ \frac{1}{3}\left(\frac{m}{|s|}+\frac{m}{|t|}+\frac{m-t}{m}\right)&\text{otherwise}\end{array}\right.$ (5)
@@
Figure 5 illustrates the Jaro similarity mean (five runs) of the subsequently generated algorithm names (i.e., $\text{Jaro}(\text{name}(a^{\prime}),\text{name}(a_{t+1}))$ for $t=0,\ldots,T-1$, where $a^{\prime}=a_{b}$ for $(1+1)$-selection and $a^{\prime}=a_{t}$ for $(1,1)$-selection) over iterations $t$ for each of the nine LLM-based optimization run configurations. In Figure 6, we show the Jaro similarity for just a single run selected for three of the nine configurations.
@@ VI Analysis of Best Algorithms
The experiment above resulted in 3 65736573\,6573 657 algorithms (out of a maximum2225555 runs with $T=100$ iterations, each, for $9$ different models. of 4 50045004\,5004 500) that were at least able to get an AOCC score larger than zero on the BBOB suite of optimization problems in $5$ dimensions. In the next step to validate our proposed framework, we evaluate how much the best of these algorithms can generalize beyond the evaluation performed during the optimization loop, and subsequently we analyze in-depth the code and behaviour of the best-performing algorithm that beats the state-the-art baselines. The evaluation is done in $5$, $10$ and $20$ dimensions, using $5$ different instances and $5$ independent runs per instance (so $25$ runs per BBOB function to optimize). For a fair comparison, we decided to re-run the generated algorithms for $d=5$ in this setting (with $25$ runs), too. The state-of-the-art baselines are the Covariance Matrix Adaptation Evolution Strategy (CMA-ES) [49], Differential Evolution (DE) [50] and an extensively optimized version of Modular CMA-ES [51, 45] denoted as CMA-best. CMA-best is a CMA-ES algorithm with active update, a Gaussian base sampler, $\lambda=20,\mu=10$, IPOP and mirrored sampling strategy and CSA step-size-adaptation (see [51] for the details of these configurations). The optimized CMA-ES algorithm is the best-configured algorithm (in terms of AOCC) on BBOB in $5$ dimensions and for a budget of 10 0001000010\,00010 000 function evaluations, out of more than 52 1285212852\,12852 128 CMA-ES and DE algorithm variants. Beating this optimized baseline is incredibly hard as it was the result of a very large modular optimization experiment. We use the CMA-best baseline to obtain a reasonable upper-bound of what is possible when optimizing an algorithm configuration to a specific benchmark set and fixed budget and dimensionality.
@@
Figure 7: The empirical attainment functions (EAF) estimate the fraction of runs that attain an arbitrary target value not later than a given runtime. The fraction ($y$-axis) denotes the cumulative fraction of target values that have been reached by an optimization algorithm, as a function of the number of function evaluations ($x$-axis). We show the EAF for the best algorithms per LLM configuration (i.e., nine algorithms) and the three baseline algorithms (CMA-ES, best CMA-ES, Differential Evolution DE), averaged over all 24 BBOB functions in $5d$.
In Figure 7 the empirical attainment functions (EAF) [47] are shown for the best algorithm generated per LLM and strategy, resulting in $9$ automatically generated algorithms plus the $3$ baselines (CMA-ES, CMA-ES-best, DE). The empirical attainment function estimates the fraction of runs that attain an arbitrary target value not later than a given runtime (i.e., number of function evaluations). Taking the partial integral of the EAF results in a more accurate version of the Empirical Cumulative Distribution Function, since it does not rely on discretization of the targets. The ”fraction” ($y$-axis in Figure 7) denotes the cumulative fraction of target values that have been reached by an optimization algorithm, as a function of the number of function evaluations ($x$-axis in Figure 7).
The area under this curve is a similar measure as the AOCC used during the LLaMEA optimization runs, and denotes how well an optimization algorithm performed over the complete run (on average over all instances, independent runs and objective functions). From the EAF curves and the area under these curves (see Table I), we can observe that LLaMEA has found one algorithm (ERADS_QuantumFluxUltraRefined, for short ERADS) with better performance for $d=5$ than the state-of-the-art CMA-ES in terms of AOCC, two algorithms (ERADS and AdaptiveDifferentialEvolutionHarmonySearch) that on average perform better than the optimized CMA-best after the total budget in $5d$ (the EAF curve reaches higher), two other algorithms that perform better than CMA-best in the first $1000$ evaluations (EnhancedFireworkAlgorithmWithLocalSearch and QuantumDifferentialParticleOptizerWithElitism) and seven algorithms that perform better than the DE baseline but worse than the CMA baseline. Detailed convergence curves per BBOB function (for $d=5$) are shown in Figure 8. From this Figure we can observe that especially for BBOB $f_{17}$ and $f_{18}$, the ERADS algorithm shows very promising search behaviour.
Figure 8: Median best-so-far function value ($y$-axis) over the 10 0001000010\,00010 000 function evaluations ($x$-axis) per BBOB function in $5d$ for the best algorithm per LLM configuration (9 algorithms: combinations of (1,1) with details, without details, and the (1+1)-variant, each with three different LLMs) and the baselines: CMA-ES, DE and CMA-best, for each of the 24 BBOB test functions.
TABLE I: Area under the EAF curve scores (AUC) for the best 3 algorithms per model in $5d$, higher AUC scores stands for a better anytime performance and have a similar meaning as the AOCC metric used earlier.
ID AUC
CMA-best (optimized baseline) 0.742
ERADS_QuantumFluxUltraRefined 0.733
CMA (baseline) 0.703
AdaptiveHybridCMAESDE 0.697
QuantumDifferentialParticleOptimizerWithElitism 0.695
EnhancedFireworkAlgorithmWithLocalSearch 0.684
AdaptiveHybridDEPSOWithDynamicRestart 0.684
ADEM 0.667
AdaptiveDifferentialEvolutionHarmonySearch 0.643
EnhancedDynamicPrecisionBalancedEvolution 0.641
DE (baseline) 0.628
QPSO 0.604
@@
38:return $\mathbf{x}_{opt},f_{opt}$triangleright Return best solution and quality
The main differences between ERADS and DE are the use of a memory factor to guide the mutation in certain directions and an adaptive $F$ and $CR$ strategy. It is unclear, why the LLM generated the ”Quantum Flux” component in the algorithm’s name, but it can be assumed that LLM hallucinations [52] will also occur when applied to code generation tasks.
@@
The key differences primarily revolve around the use of a memory vector and a memory factor in combination with parameter adaptation. JADE uses an optional archive with the best solutions which primarily serves to maintain diversity. Furthermore, JADE also features adaptive control parameters ($F$ and $CR$) that adjust dynamically based on the success rates of previous generations, aiming to optimise these parameters in real time to improve performance. It is interesting to observe that most of the best algorithms proposed by the LLM are similar to Differential Evolution (or Hybrids of CMA and DE), which could indicate a bias towards this kind of algorithms in the LLM, or it could indicate that overall these kind of algorithms work well for the specific benchmark suite.